# Quantum Search with a Generalized Laplacian

### v1

@@ II Spin Networks and Signed Weighted Graphs
[…]
@@
=($=\left(\sum_{\{i,j\}}e_{ij}-2e_{12}\right){\left|1\right\rangle}$
@@
$H_{Z}{\left|2\right\rangle}$ =($=\left(\sum_{\{i,j\}}e_{ij}-2e_{21}-2e_{23}-2e_{24}\right){\left|2\right\rangle}$
@@
$H_{Z}{\left|3\right\rangle}$ =($=\left(\sum_{\{i,j\}}e_{ij}-2e_{32}-2e_{34}\right){\left|3\right\rangle}$
@@
$H_{Z}{\left|2\right\rangle}$ =($=\left(\sum_{\{i,j\}}e_{ij}-2e_{42}-2e_{43}\right){\left|4\right\rangle}$
@@
$H_{Z}{\left|k\right\rangle}$ =($=\left(\sum_{\{i,j\}}e_{ij}-2\sum_{\ell\sim k}e_{k\ell}\right){\left|k\right\rangle}$
@@ III.1 Algorithm
$Z_{a}{\left|i\right\rangle}=\begin{cases}{\left|i\right\rangle},&i=a,\\ -{\left|i\right\rangle},&i\neq a,\\ \end{cases}$
@@
$-\frac{1}{2}Z_{a}=-{\left|a\middle\rangle\middle\langle a\right|},$
@@ III.2 Numerical Simulations
Figure 3: Success probability vs time for search on the weighted barbell graph with $N=1200$ vertices using the generalized Laplacian quantum walk with (a) $\alpha=-3$, (b) $\alpha=0$, (c) $\alpha=1$, (d) $\alpha=2$, and (e) $\alpha=4$. In all the left plots, the solid black curve is when the weight of the bridge is $w=120$, dashed red is $w=150$, dotted green is $w=200$, dot-dashed blue is $w=300$, dot-dot-dashed orange is $w=600$, and dot-dashed-dashed brown is $w=1200$. In all the right plots, the respective curves are when the weights are negated, i.e., $w=-120$, $-150$, $-200$, $-300$, $-600$, and $-1200$.
In Fig. 3, we plot the success probability of the algorithm as it evolves with time for search on the weighted barbell graph of $N=1200$ vertices. Figure 3a uses the the generalized Laplacian with $\alpha=-3$, and it consists of two plots: a left plot with various positive weights for the bridge, $w=120$, $150$, $200$, $300$, $600$, and $1200$, and a right plot with the negatives of these weights, i.e., $w=-120$, $-150$, $-200$, $-300$, $-600$, and $-1200$. Most of the curves overlap, reaching a success probability of $1/2$ at time $(\pi/2)\sqrt{N/2}\approx 38.5$ because the graph acts like two disconnected cliques. There are two critical weights with negative values, however, $w=-120$ and $w=-200$, where probability is able to move across the bridge, and a higher success probability is obtained.
In Fig. 3b, the success probability is plotted for search with the generalized Laplacian with $\alpha=0$, i.e., with the standard Laplacian. The left plot shows that positive weights do not help probability move across the bridge, as proved in [12]. The right plot, however, shows that there is a negative critical weight, $w=-300$, where the success probability is improved, and this was not discovered in [12] because that work only considered positive weights.
The generalized Laplacian with $\alpha=1$, which is the adjacency matrix, was used in Fig. 3c. In the left plot, there is a positive critical weight $w=600$ that boosts the success probability, and it was proved in [12] that this corresponds to a weight of $w=N/2$ and reaches a success probability of $0.820$ (independent of $N$) at a time of $87.22.518\sqrt{N}=$ 87.2. Ref. [12] did not consider negative weights, but we do in the right plot, and we see that there is a negative critical weight of $w=-600$ also boosts the success probability.
@@
Note $\alpha$ need not be an integer, and additional simulations show that the $\alpha<0$ cases all behave similarly: there are two negative critical weights with boosted success probability. The $\alpha=0$ case (Laplacian) has one negative critical weight. When $10<\alpha<$ 1, there is one positive critical weight and one negative critical weight. When $\alpha=1$ (adjacency), there is one positive critical weight. When $\alpha>1$, there are two positive critical weights. Furthermore, for the weighted barbell graph with $N=1200$ vertices, the boosted success probability either occurs at time $87.2$ or at time $208.2$. Next, we will analytically derive these results, beginning with the critical weights for general $N$ and $\alpha$, followed by the success probabilities and runtimes.
@@ III.3 Noncritical and Critical Weights
${\left|c\right\rangle},\quad{\left|d\right\rangle},\quad{\left|e\right\rangle}=\frac{1}{\sqrt{N/2-2}}\sum_{i\in e}{\left|i\right\rangle}.$
@@
These two critical weights are computed in Table 1 for a weighted barbell graph of $N=1200$ vertices for various parameters $\alpha$ of the generalized Laplacian. These are in agreement with Fig. 3, which showed that $\alpha=-3$ had critical weights of $-200$ and $-120$, $\alpha=0$ had one critical weight of $-300$, $\alpha_{1}$ had two critical weights of $600$ and $-600$, $\alpha=2$ had one critical weight of $300$, and $\alpha=4$ had two critical weights of $150$ and $300$. The formulas (11) and (12) also reveal why there are two negative critical weights when $\alpha<0$, one negative critical weight when $\alpha=0$, one positive and one negative critical weight when $10<\alpha<$ 1, one positive critical weight when $\alpha=1$, and two positive critical weights when $\alpha>1$. It also reveals why we get one evolution when $w=N/(2\alpha)$ and another when $w=N/[2(\alpha-2)]$, i.e., either ${\left|cd_{+}\right\rangle}$ or ${\left|cd_{-}\right\rangle}$ is degenerate with ${\left|a\right\rangle}$, ${\left|b\right\rangle}$, and ${\left|e\right\rangle}$, respectively. The other weights in the plots are noncritical weights, where probability cannot appreciably cross the bridge.
@@ III.4 Behavior when $w$ is noncritical
$=\frac{e^{it}}{2\sqrt{2}}\left[\left(e^{i\sqrt{2/N})t}-e^{-i\sqrt{2/N})t}\right){\left|a\right\rangle}\right.$
$\quad\quad\quad\enspace\left.+\left(e^{i\sqrt{2/N})t}+e^{-i\sqrt{2/N})t}\right){\left|b\right\rangle}\right]+\frac{1}{\sqrt{2}}e^{it}{\left|e\right\rangle}$
@@ III.6 Behavior when $w=w_{-}$
Note that the success probability $p_{a}(t)$ has strong agreement with Fig. 3 when $N=1200$ and $w=w_{-}$. To find the time at which the success probability peaks, we set the derivative of $p_{a}(t)$ equal to 00 and solve for $t$. The derivative of the success probability is
@@
×$\quad\times\left[\sqrt{2-\sqrt{2}}\sqrt{\frac{2+\sqrt{2}}{N}}\cos\left(\sqrt{\frac{2+\sqrt{2}}{N}}t\right)+\sqrt{2+\sqrt{2}}\sqrt{\frac{2-\sqrt{2}}{N}}\cos\left(\sqrt{\frac{2-\sqrt{2}}{N}}t\right)\right]$
@@
×$\quad\times\left[\cos\left(\sqrt{\frac{2+\sqrt{2}}{N}}t\right)+\cos\left(\sqrt{\frac{2-\sqrt{2}}{N}}t\right)\right].$
@@
For example, when $N=1200$, the runtime is $t_{*}\approx 208.2$, in strong agreement with Fig. 3 when $w=w_{-}$. Plugging $t_{*}$ into $p_{a}(t)$ and letting $A=\sqrt{2+\sqrt{2}}$ and $B=\sqrt{2-\sqrt{2}}$, the success probability reaches a height of
@@ Appendix A Two-stage algorithm when $w=w_{+}$ maximizing the probability in $a$ and $b$
$p_{ab}(t)$ $=\frac{1}{8}\left\{\left[\sqrt{2+\sqrt{2}}\sin\left(\sqrt{\frac{2+\sqrt{2}}{N}}t\right)\right.\right.$
@@
×$\qquad\enspace\times\left[\left(1+\sqrt{2}\right)\cos\left(\sqrt{\frac{2+\sqrt{2}}{N}}t\right)\right.$
@@ Appendix B Two-stage algorithm when $w=w_{-}$ maximizing the probability in the marked clique
This is consistent with Fig. 6a, where around time $211.26.097\sqrt{1200}=$ 211.2, the probability in the marked clique reaches $0.917$. The state of the system at the end of the first stage is ${\left|\psi(t_{1})\right\rangle}$ and can be obtained from (15), and we drop the overall phase of $e^{it}$ to obtain
@@
Since $p_{a}(t)$ is a squared sinusoidal function with amplitude $0.840$, during the second stage of the algorithm, the success probability reaches a maximum of $0.840$, as reported in the main text. The time that this takes is the value of $\Delta t$, call it $t_{2}$, that makes the argument of the sine equal to $23\pi/$ 2, i.e.,

### v2

@@ II Spin Networks and Signed Weighted Graphs
A spin-1/2 particle has two eigenstates in $Z$, denoted spin up $\left|\uparrow\right\rangle$} and spin down $\left|\downarrow\right\rangle$}, with corresponding eigenvalues $1$ and 1-1. The Pauli matrices act on these eigenstates of $Z$ by
@@
$=\left(\sum_{\{i,j\}}e_{ij}-2e_{12}\right){\left|1\right\rangle}$
@@
$H_{Z}{\left|2\right\rangle}$ $=\left(\sum_{\{i,j\}}e_{ij}-2e_{21}-2e_{23}-2e_{24}\right){\left|2\right\rangle}$
@@
$H_{Z}{\left|3\right\rangle}$ $=\left(\sum_{\{i,j\}}e_{ij}-2e_{32}-2e_{34}\right){\left|3\right\rangle}$
@@
$H_{Z}{\left|2\right\rangle}$ $=\left(\sum_{\{i,j\}}e_{ij}-2e_{42}-2e_{43}\right){\left|4\right\rangle}$
@@
$H_{Z}{\left|k\right\rangle}$ $=\left(\sum_{\{i,j\}}e_{ij}-2\sum_{\ell\sim k}e_{k\ell}\right){\left|k\right\rangle}$
@@ III.1 Algorithm
[…]
@@
$\frac{1}{2}Z_{a}=-{\left|a\middle\rangle\middle\langle a\right|},$
@@ III.2 Numerical Simulations
Figure 3: Success probability vs time for search on the weighted barbell graph with $N=1200$ vertices using the generalized Laplacian quantum walk with (a) $\alpha=-3$, (b) $\alpha=0$, (c) $\alpha=1$, (d) $\alpha=2$, and (e) $\alpha=4$. In all the left plots, the solid black curve is when the weight of the bridge is $w=120$, dashed red is $w=150$, dotted green is $w=200$, dot-dashed blue is $w=300$, dot-dot-dashed orange is $w=600$, and dot-dashed-dashed brown is $w=1200$. In all the right plots, the respective curves are when the weights are negated, i.e., w=120w=-120, 150-150, 200-200, 300-300, 600-600, and 1200-1200.
In Fig. 3, we plot the success probability of the algorithm as it evolves with time for search on the weighted barbell graph of $N=1200$ vertices. Figure 3a uses the the generalized Laplacian with $\alpha=-3$, and it consists of two plots: a left plot with various positive weights for the bridge, $w=120$, $150$, $200$, $300$, $600$, and $1200$, and a right plot with the negatives of these weights, i.e., w=120w=-120, 150-150, 200-200, 300-300, 600-600, and 1200-1200. Most of the curves overlap, reaching a success probability of $1/2$ at time $(\pi/2)\sqrt{N/2}\approx 38.5$ because the graph acts like two disconnected cliques. There are two critical weights with negative values, however, w=120w=-120 and w=200w=-200, where probability is able to move across the bridge, and a higher success probability is obtained.
In Fig. 3b, the success probability is plotted for search with the generalized Laplacian with $\alpha=0$, i.e., with the standard Laplacian. The left plot shows that positive weights do not help probability move across the bridge, as proved in [12]. The right plot, however, shows that there is a negative critical weight, w=300w=-300, where the success probability is improved, and this was not discovered in [12] because that work only considered positive weights.
The generalized Laplacian with $\alpha=1$, which is the adjacency matrix, was used in Fig. 3c. In the left plot, there is a positive critical weight $w=600$ that boosts the success probability, and it was proved in [12] that this corresponds to a weight of $w=N/2$ and reaches a success probability of $0.820$ (independent of $N$) at a time of $2.518\sqrt{N}=87.2$. Ref. [12] did not consider negative weights, but we do in the right plot, and we see that there is a negative critical weight of w=600w=-600 also boosts the success probability.
@@
Note $\alpha$ need not be an integer, and additional simulations show that the $\alpha<0$ cases all behave similarly: there are two negative critical weights with boosted success probability. The $\alpha=0$ case (Laplacian) has one negative critical weight. When $0<\alpha<2$, there is one positive critical weight and one negative critical weight. When $\alpha=2$ (signless Laplacian), there is one positive critical weight. When $\alpha>2$, there are two positive critical weights. Furthermore, for the weighted barbell graph with $N=1200$ vertices, the boosted success probability either occurs at time $87.2$ or at time $208.2$. Next, we will analytically derive these results, beginning with the critical weights for general $N$ and $\alpha$, followed by the success probabilities and runtimes.
@@ III.3 Noncritical and Critical Weights
${\left|c\right\rangle},\quad{\left|d\right\rangle},\quad{\left|e\right\rangle}=\frac{1}{\sqrt{N/2-1}}\sum_{i\in e}{\left|i\right\rangle}.$
@@
These two critical weights are computed in Table 1 for a weighted barbell graph of $N=1200$ vertices for various parameters $\alpha$ of the generalized Laplacian. These are in agreement with Fig. 3, which showed that $\alpha=-3$ had critical weights of 200-200 and 120-120, $\alpha=0$ had one critical weight of 300-300, $\alpha=1$ had two critical weights of $600$ and 600-600, $\alpha=2$ had one critical weight of $300$, and $\alpha=4$ had two critical weights of $150$ and $300$. The formulas (11) and (12) also reveal why there are two negative critical weights when $\alpha<0$, one negative critical weight when $\alpha=0$, one positive and one negative critical weight when $0<\alpha<2$, one positive critical weight when $\alpha=2$, and two positive critical weights when $\alpha>2$. It also reveals why we get one evolution when $w=N/(2\alpha)$ and another when $w=N/[2(\alpha-2)]$, i.e., either $\left|cd_{+}\right\rangle$} or $\left|cd_{-}\right\rangle$} is degenerate with $\left|a\right\rangle$}, $\left|b\right\rangle$}, and $\left|e\right\rangle$}, respectively. The other weights in the plots are noncritical weights, where probability cannot appreciably cross the bridge.
@@ III.4 Behavior when $w$ is noncritical
$=\frac{e^{it}}{2\sqrt{2}}\left[\left(e^{i\sqrt{2/N}t}-e^{-i\sqrt{2/N}t}\right){\left|a\right\rangle}\right$.
$\quad\quad\quad\enspace\left.+\left(e^{i\sqrt{2/N}t}+e^{-i\sqrt{2/N}t}\right){\left|b\right\rangle}\right]+\frac{1}{\sqrt{2}}e^{it}{\left|e\right\rangle}$
@@ III.6 Behavior when $w=w_{-}$
Note that the success probability $p_{a}(t)$ has strong agreement with Fig. 3 when $N=1200$ and $w=w_{-}$. To find the time at which the success probability peaks, we set the derivative of $p_{a}(t)$ equal to 0 and solve for $t$. The derivative of the success probability is
@@
$\quad\times\left[\sqrt{2-\sqrt{2}}\sqrt{\frac{2+\sqrt{2}}{N}}\cos\left(\sqrt{\frac{2+\sqrt{2}}{N}}t\right)+\sqrt{2+\sqrt{2}}\sqrt{\frac{2-\sqrt{2}}{N}}\cos\left(\sqrt{\frac{2-\sqrt{2}}{N}}t\right)\right]$
@@
$\quad\times\left[\cos\left(\sqrt{\frac{2+\sqrt{2}}{N}}t\right)+\cos\left(\sqrt{\frac{2-\sqrt{2}}{N}}t\right)\right].$
@@
For example, when $N=1200$, the runtime is $t_{*}\approx 208.2$, in strong agreement with Fig. 3 when $w=w_{-}$. Note $k=1,3$ respectively yield the local maxima at $t=41.6,124.9$ in Fig. 3 when $w=w_{-}$, which is why $k=5$ was chosen in (17). Plugging $t_{*}$ into $p_{a}(t)$ and letting $A=\sqrt{2+\sqrt{2}}$ and $B=\sqrt{2-\sqrt{2}}$, the success probability reaches a height of
@@ Appendix A Two-stage algorithm when $w=w_{+}$ maximizing the probability in $a$ and $b$
[…]
@@
$\qquad\enspace\times\left[\left(1+\sqrt{2}\right)\cos\left(\sqrt{\frac{2+\sqrt{2}}{N}}t\right)\right$.
@@ Appendix B Two-stage algorithm when $w=w_{-}$ maximizing the probability in the marked clique
This is consistent with Fig. 6a, where around time $6.097\sqrt{1200}=211.2$, the probability in the marked clique reaches $0.917$. The state of the system at the end of the first stage is $\left|\psi(t_{1})\right\rangle$} and can be obtained from (15), and we drop the overall phase of $e^{it}$ to obtain
@@
Since $p_{a}(t)$ is a squared sinusoidal function with amplitude $0.840$, during the second stage of the algorithm, the success probability reaches a maximum of $0.840$, as reported in the main text. The time that this takes is the value of $\Delta t$, call it $t_{2}$, that makes the argument of the sine equal to $3\pi/2$, i.e.,